# Model Inversion Attacks Meet Cryptographic Fuzzy Extractors

### v1

@@ 4.1 Existing Scheme via Fuzzy Extractor
A fuzzy extractor (FE) is an existing cryptographic primitive that derives high-entropy cryptographic keys from a noisy input source, such as human biometrics [25]. Specifically, it extracts a uniformly random string from an input vector, such that the same string can be reliably reproduced from other close input vectors. This “error-correcting” property makes FE well-suited for noisy-source authentication.
@@
A FE primitive is defined by two functions Gen and Rep. The function Gen takes input $x$ and outputs a secret extracted key $r$ and a public helper string $p$. The function Rep takes some close input $x^{\prime}$ and helper $p$ to reproduce $r$. Typically, an FE scheme is instantiated using another cryptographic primitive called a secure sketch, which provides the error-correction property, and a randomness extractor, which ensures uniformity of the secret $r$ [25]. The security property of FE guarantees that the helper $p$ can be released publicly without revealing any information about $r$, enabling $r$ to be used as a secret key for any downstream cryptographic applications.
@@
The roles of the FE Gen outputs $r$ and $p$ as secrets and helpers respectively correspond to the outputs $y$ and $s$ of the ideal primitive we defined earlier. Hence, an FE scheme can be easily integrated into the face authentication pipeline in Algo. 1 by substituting with FE’s Gen and Rep functions  [^3] Other methods of integrating FE into face authentication have also been proposed, such as [33] using the FE secret output $r$ to generate keys for a public-key encryption scheme, though this distinction is not relevant.. Note that only $(r,p)$ must be persistently stored for subsequent use at authentication time.
No secure FE construction prior to our work that performs $\ell_{2}$ norm error-correction is known to the best of our knowledge. To illustrate the lack of such a primitive, we will take the closest FE scheme we know of that approximates for $\ell_{2}$-correction, called Facial-Fuzzy Extractor (or Facial-FE) [23], and show that it is insecure.
@@
Facial-FE claims to provide the only computationally secure FE that works for an approximation of the $\ell_{2}$ metric using a specific lattice-based metric [23]. It uses dense and efficiently decodable lattices, where each input is decoded to a lattice point, and the difference vector is the helper string. The hashed input and the public helper string are stored as the persistent secret to help reconstruct the original input from the authentication input. The Facial-FE construction from [23] using the E8-lattice decoder as the Decode sub-algorithm is reproduced as-is in Algo. 2, adapted to use symbols consistent with our work. $H_{k}$ denotes a cryptographic hash with a random key $k$.
@@
Facial-FE can satisfy Item 1 because the linearity of lattices allows for correction of errors within the Voronoi cell, which is an approximation for correction of errors within some $\ell_{2}$ distance. However, it does not satisfy the security property Item 2. Notably, it releases a public helper string $p$, which contains a string $ss$ (called a secure sketch output). The key issue is that the construction uses a deterministic decoder, i.e., there is no randomization in producing $ss$. Since $c$ is determined by $x$ for a fixed lattice, $ss$ = cxc-x can reveal a lot of information about $x$. Our attacks in Section 5 take advantage of this weakness and use $ss$ to find an $x^{\prime}$ close to $x$.
@@
Algorithm 2 Facial-FE construction from [23] Algo. 1, 2, 5, 6.
@@ 4.2 Existing Scheme via Random Projection
Another popularly used primitive for protecting embeddings via post-processing is multispace random projection (MRP) [24]. There are many projection-based schemes used to protect biometrics, but they are variants of the common MRP building block [34, 35]. MRP works by projecting input vectors onto a random subspace of lower dimensionality, determined by a user-specific random seed. Then it can perform vector comparisons for authentication in the projected vector space without storing the original vectors.
@@
MRP is a randomized technique that projects the input vector along random vectors, first proposed by [24]. MRP samples a random Gaussian matrix $R\in\mathbb{R}^{n\times m}$ where each entry is from the standard normal distribution with mean 0 and standard deviation $1$. Each row vector in $R$ thus has a random vector direction in the Euclidean space. It projects the input vector $x\in\mathbb{R}^{m}$ to output vector $y\in\mathbb{R}^{n}$, where $n<m$, as follows:
@@
MRP does meet $\ell_{2}$ norm correctness Item 1 because of the phenomenon captured by the Johnson-Lindenstrauss (JL) Lemma [36, 37]. The JL Lemma 1 given below ensures that the pairwise distance between all points in a set of $k$ points is maintained within an arbitrarily small factor by random projection maps. In other words, the projected data still retains enough information for authentication to work.
Lemma 1 (Johnson–Lindenstrauss Lemma [37] rephrased).
@@ 5.1 Attack Setup
We consider the attack setup where the seed $s$ is a different value per-user. The use of different seeds is similar to how passwords of each user are hashed with different salt values, which in the context of password hashing, is expected to increase the attacker’s effort per user. It has been considered in prior works on protecting embeddings as well. For example, the MRP seed is referred to as a “user-specific PRN” [24], and the Facial-FE seed includes a user-specific deterministic vector $ss$.
@@ 5.2.1 Step 1: Recovering surrogate embedding
In the case of MRP, the attacker wants to recover some $x^{*}$ from $y=Rx$. Since $R\in\mathbb{R}^{n\times m}$ and $n<m$, $R$ is a compression map. Solving $x=R^{-1}y$ is an under-determined system of linear equations, so Gaussian elimination is not usable directly. To resolve this problem, we propose to compute a Moore-Penrose inverse [38] or simply, pseudo-inverse of $R$, denoted as $R^{\dagger}\in\mathbb{R}^{m\times n}$ using the singular values of $R$. The matrix $R^{\dagger}$ thus obtained is a right pseudo-inverse of $R$ as $m<n$. By left multiplying $R^{\dagger}$ to the leaked vector $y$, we get an $m$-dimensional vector $x^{*}$.
@@ 5.2.2 Step 2: Inverting embedding
There are many existing ML models that can be considered for use as inversion models. Most create good quality images but dissimilar images [12], or poor quality but representative images [39, 14]. Moreover, these models have only been examined in the setting of image inversion from “unprotected” embeddings $x$ and not from approximately recovered $x^{*}$ as needed in our attack. For our inversion model, we adapt the classifier-guided diffusion model by [40] and use surrogate vectors $x^{*}$ obtained in Step 1 instead of class labels as the guidance vector. The public face datasets are used to train the model to learn an association between the distribution of face images $I$ and surrogate embeddings $x^{*}$. Once trained, the inversion model is fixed for use in attacks. At attack time (post-training), the inversion model then generates images by conditioning on the surrogate $x^{*}$ recovered by Step 1 from the leaked embedding $x$.
@@ 6.1 Methodology & Attack Setup
Recall that Pipe uses an embedding conditioned diffusion model adapted from the implementation by [40] as the inversion model in Pipe attack. We use hyperparameters and image size of $64\times 64$ as recommended in [40], with a learning rate of 1e41e-4, batch_size =64, ema_rate of $0.9999$ and anneal_steps 25000.
@@
To illustrate that our attack can work across multiple datasets and face distributions, we use $3$ image datasets: Casia-Webface [43], Celeb-A [45], LFW [47]. We perform cross-dataset evaluation. Specifically, the inversion model used in Pipe is trained only on the Celeb-A dataset, while the ASR is reported for all the $3$ datasets separately. Furthermore, the inversion model is tested on $1000$ different users (images) in all $3$ datasets, but trained only on $9177$ user images in the Celeb-A dataset different from the ones tested on. Thus, there is no overlap in identities used in training set of Pipe and those tested.
@@ 6.2 PIPE Against Existing Protection Schemes
Table III shows that Pipe succeeds with a high ASR. When setting up Facial-FE, we configured its parameters to attain $\leq 70\%$ FNR [^5] TPR is 11-FNR. FNR $\leq 70\%$ implies TPR $30\%$. as specified by [23]. This requires the face vectors to be scaled down. The scales we obtain for Facenet and Arcface embeddings are 5.7 and 1.39 respectively. At this scale, which is required for reasonable accuracy, we notice that the inputs tend to decode to lattice points $c$ that are sparse 512-dimensional vectors with more than 194 zero elements. Consequently, the helper string ss=cxss=c-x contains significant information about $x$ making Facial-FE insecure . This issue cannot be resolved simply by changing the scale factor. As mentioned in [23] as well, when we experimentally reduce the scale factor to improve security significantly, the TPR becomes unusably small. Thus, the poor trade-off between security and accuracy appears unavoidable for Facial-FE.
@@ 6.3 Prior Attacks vs Partial Protection Schemes
To answer this, we evaluate $3$ state-of-the-art inversion attacks on face recognition models: Bob [14], KED-MI [30], GMI [31]. Bob is the state-of-the-art white-box embedding inversion attack which has reported better ASR than attacks preceding it. KED-MI and GMI are state-of-the-art black-box model inversion attacks [^6] The core of these attacks only requires black-box access to the target model, and a separate inversion model trained on the public dataset.  [17]. They both use GANs as the inversion model and have the best-known performance among black-box model inversion attacks [15].
@@ Results.
Figure 3 presents the representative images reconstructed by Pipe and Bob. Most Pipe reconstructions are accepted as the same face by the target model, which is not the case for Bob. Moreover, as seen in Tables VIII and IX, Pipe achieves a lower (better) visual similarity LPIPS score [9] and a lower (better) FID score [48]. This further confirms that Pipe creates perceptually closer reconstructions to the original images (see Appendix E for details).
KED-MI and GMI attacks obtain less than $0.16\%$ ASR on average over 5 trials, with or without a protection scheme. The weak performance of KED-MI and GMI is expected since these attacks rely solely on classification outputs, which are inherently more restrictive than using the embedding vector itself. We qualitatively inspected over $20$ examples returned by these attacks, and we find that the reconstructions are quite far from the original consistently. Our implementation builds on code provided by [15].
@@ 7 Our Proposed Scheme: L2FE-Hash
The Rep procedure receives $x^{\prime}$. It de-serializes $p$ to obtain $c=Ab+x$ that was computed during Gen. It implicitly checks if $c-x^{\prime}$ decodes to the correct ephemeral secret value by checking if the hash of the decoded value $H_{k}(b)$ matches what is stored in the database. The Decode procedure is the well-known Babai’s Nearest Plane (BNP) algorithm [49], which performs the $\ell_{2}$ error correction.
@@
Constructing a fuzzy extractor for an arbitrary family of distributions $\mathcal{X}$ is known to be statistically impossible [50]. Therefore, to characterize the computational hardness of breaking L2FE-Hash, one must specialize to specific distributions. If we strictly confine the input face distribution $\cal{X}$ to be discrete Gaussian or uniform over a small interval, then the hardness of recovering $b$ from $(A,Ab+x)$ is exactly the cryptographic hardness of the LWE problem [51, 52]. Nonetheless, this simplification is too unrealistic, as real-world face distributions are not known to follow any cryptographically hard distribution or have any closed-form mathematical characterization.
As the next best approach, we characterize the support of the distribution $\mathcal{X}$ as a set of balls bounded in some $R,R]^{m}$ space, since our application aims to correct within some $\ell_{2}$ radius for a set of users. We refer to such distributions with high min-entropy $h$ as well-behaved. Under the assumption that $\mathcal{X}$ is uniform over the assumed support, we prove that L2FE-Hash is a secure FE where the number of secure output bits $\kappa$ is upper bounded by some function of $h$. Theorem 1 makes the claim precise (see Appendix B for the proof). The L2FE-Hash output $r$ is a $\kappa$-bitstring, but equivalently a vector in $\mathbb{Z}_{q}^{n}$ as needed, where $n=\kappa\log_{q}(2)$.
@@
Theorem 1 specifies the minimum input entropy needed for good security for an abstract characterization of the distribution $\cal{X}$. But, determining entropy $h$ for real distributions is intractable. So, we derive practical parameters such that the protected embeddings give good accuracy (TPR over $60\%$) and tractable runtime for Gen and Rep. Note that once TPR exceeds well over $50\%$, majority voting over multiple independent face samples can boost it arbitrarily close to $100\%$ [53]. To achieve accuracy of $1-\delta$, it suffices to have $O(\log(1/\delta))$ face images of the user at authentication time, which can be run in parallel for efficiency. We evaluate L2FE-Hash using the best-known attacks, including prior ones and our Pipe attack on real datasets, to practically assess security.
@@
We have shown that L2FE-Hash is a provably secure fuzzy extractor for $\ell_{2}$ distance if the input distribution $\cal{X}$ satisfies certain conditions. Next, we state the desired connection between a fuzzy extractor and ideal defense primitive as Theorem 2. We defer the formal definition of these two primitives and the ideal primitive security game to Appendix A, and the proof of Theorem 2 to Appendix C due to space limits.
@@
Let FE be a ($X,\mathcal{X},\kappa,t,\epsilon_{\text{FE}}$)-fuzzy extractor as in [25]. Then FE is a ($X,\mathcal{X},\{0,1\}^{\kappa},t,\lambda$)-ideal primitive with adversarial advantage $\text{Adv}(\mathcal{A})\leq\epsilon_{\text{FE}}(1-\delta)$.

### v2

@@ 4.1 Existing Scheme via Fuzzy Extractor
A fuzzy extractor (FE) is an existing cryptographic primitive that derives high-entropy cryptographic keys from a noisy input source, such as human biometrics [2004:fe]. Specifically, it extracts a uniformly random string from an input vector, such that the same string can be reliably reproduced from other close input vectors. This “error-correcting” property makes FE well-suited for noisy-source authentication.
@@
A FE primitive is defined by two functions Gen and Rep. The function Gen takes input $x$ and outputs a secret extracted key $r$ and a public helper string $p$. The function Rep takes some close input $x^{\prime}$ and helper $p$ to reproduce $r$. Typically, an FE scheme is instantiated using another cryptographic primitive called a secure sketch, which provides the error-correction property, and a randomness extractor, which ensures uniformity of the secret $r$ [2004:fe]. The security property of FE guarantees that the helper $p$ can be released publicly without revealing any information about $r$, enabling $r$ to be used as a secret key for any downstream cryptographic applications.
@@
The roles of the FE Gen outputs $r$ and $p$ as secrets and helpers respectively correspond to the outputs $y$ and $s$ of the ideal primitive we defined earlier. Hence, an FE scheme can be easily integrated into the face authentication pipeline in Algo. 1 by substituting with FE’s Gen and Rep functions  [^3] Other methods of integrating FE into face authentication have also been proposed, such as [2017:fe-pke] using the FE secret output $r$ to generate keys for a public-key encryption scheme, though this distinction is not relevant.. Note that only $(r,p)$ must be persistently stored for subsequent use at authentication time.
No secure FE construction prior to our work that performs $\ell_{2}$ norm error-correction is known to the best of our knowledge. To illustrate the lack of such a primitive, we will take the closest FE scheme we know of that approximates for $\ell_{2}$-correction, called Facial-Fuzzy Extractor (or Facial-FE) [2021:facial-fe], and show that it is insecure.
@@
Facial-FE claims to provide the only computationally secure FE that works for an approximation of the $\ell_{2}$ metric using a specific lattice-based metric [2021:facial-fe]. It uses dense and efficiently decodable lattices, where each input is decoded to a lattice point, and the difference vector is the helper string. The hashed input and the public helper string are stored as the persistent secret to help reconstruct the original input from the authentication input. The Facial-FE construction from [2021:facial-fe] using the E8-lattice decoder as the Decode sub-algorithm is reproduced as-is in Algo. 2, adapted to use symbols consistent with our work. $H_{k}$ denotes a cryptographic hash with a random key $k$.
@@
Facial-FE can satisfy Item 1 because the linearity of lattices allows for correction of errors within the Voronoi cell, which is an approximation for correction of errors within some $\ell_{2}$ distance. However, it does not satisfy the security property Item 2. Notably, it releases a public helper string $p$, which contains a string $ss$ (called a secure sketch output). The key issue is that the construction uses a deterministic decoder, i.e., there is no randomization in producing $ss$. Since $c$ is determined by $x$ for a fixed lattice, $ss$ = cxc-x can reveal a lot of information about in $x$. Our attacks in Section 5 take advantage of this weakness and use $ss$ to find an $x^{\prime}$ close to $x$.
@@
Algorithm 2 Facial-FE construction from [2021:facial-fe] Algo. 1, 2, 5, 6.
@@ 4.2 Existing Scheme via Random Projection
Another popularly used primitive for protecting embeddings via post-processing is multispace random projection (MRP) [2007:mrp]. There are many projection-based schemes used to protect biometrics, but they are variants of the common MRP building block [2004:biohashing, 2018:iom-hash]. MRP works by projecting input vectors onto a random subspace of lower dimensionality, determined by a user-specific random seed. Then it can perform vector comparisons for authentication in the projected vector space without storing the original vectors.
@@
MRP is a randomized technique that projects the input vector along random vectors, first proposed by [2007:mrp]. MRP samples a random Gaussian matrix $R\in\mathbb{R}^{n\times m}$ where each entry is from the standard normal distribution with mean 0 and standard deviation $1$. Each row vector in $R$ thus has a random vector direction in the Euclidean space. It projects the input vector $x\in\mathbb{R}^{m}$ to output vector $y\in\mathbb{R}^{n}$, where $n<m$, as follows:
@@
MRP does meet $\ell_{2}$ norm correctness Item 1 because of the phenomenon captured by the Johnson-Lindenstrauss (JL) Lemma [1984:JLLemma, 2003:JLLemma-proof]. The JL Lemma 1 given below ensures that the pairwise distance between all points in a set of $k$ points is maintained within an arbitrarily small factor by random projection maps. In other words, the projected data still retains enough information for authentication to work.
Lemma 1 (Johnson–Lindenstrauss Lemma [2003:JLLemma-proof] rephrased).
@@ 5.1 Attack Setup
We consider the attack setup where the seed $s$ is a different value per-user. The use of different seeds is similar to how passwords of each user are hashed with different salt values, which in the context of password hashing, is expected to increase the attacker’s effort per user. It has been considered in prior works on protecting embeddings as well. For example, the MRP seed is referred to as a “user-specific PRN” [2007:mrp], and the Facial-FE seed includes a user-specific deterministic vector $ss$. Choosing the same seed for all users would make our attack much more effective, though that would be an artificial choice.
@@ 5.2.1 Step 1: Recovering surrogate embedding
In the case of MRP, the attacker wants to recover some $x^{*}$ from $y=Rx$. Since $R\in\mathbb{R}^{n\times m}$ and $n<m$, $R$ is a compression map. Solving $x=R^{-1}y$ is an under-determined system of linear equations, so Gaussian elimination is not usable directly. To resolve this problem, we propose to compute a Moore-Penrose inverse [1955:moore-penrose-inverse] or simply, pseudo-inverse of $R$, denoted as $R^{\dagger}\in\mathbb{R}^{m\times n}$ using the singular values of $R$. The matrix $R^{\dagger}$ thus obtained is a right pseudo-inverse of $R$ as $m<n$. By left multiplying $R^{\dagger}$ to the leaked vector $y$, we get an $m$-dimensional vector $x^{*}$.
@@ 5.2.2 Step 2: Inverting embedding
There are many existing ML models that can be considered for use as inversion models. Most create good quality images but dissimilar images [2023:stylegan-recontruction-attack], or poor quality but representative images [2022:cnn-recontruction-attack, 2024:bob]. Moreover, these models have only been examined in the setting of image inversion from “unprotected” embeddings $x$ and not from approximately recovered $x^{*}$ as needed in our attack. For our inversion model, we adapt the classifier-guided diffusion model by [2021:conditionedDDPM] and use surrogate vectors $x^{*}$ obtained in Step 1 instead of class labels as the guidance vector. The public face datasets are used to train the model to learn an association between the distribution of face images $I$ and surrogate embeddings $x^{*}$. Once trained, the inversion model is fixed for use in attacks. At attack time (post-training), the inversion model then generates images by conditioning on the surrogate $x^{*}$ recovered by Step 1 from the leaked embedding $x$.
@@ 6.1 Methodology & Attack Setup
Recall that Pipe uses an embedding conditioned diffusion model adapted from the implementation by [2021:conditionedDDPM] as the inversion model in Pipe attack. We use hyperparameters and image size of $64\times 64$ as recommended in [2021:conditionedDDPM], with a learning rate of 1e41e-4, batch_size =64, ema_rate of $0.9999$ and anneal_steps 25000.
@@
To illustrate that our attack can work across multiple datasets and face distributions, we use $3$ image datasets: Casia-Webface [2014:casia-webface], Celeb-A [2015:celeba], LFW [21012:lfw]. We perform cross-dataset evaluation. Specifically, the inversion model used in Pipe is trained only on the Celeb-A dataset, while the ASR is reported for all the $3$ datasets separately. Furthermore, the inversion model is tested on $1000$ different users (images) in all $3$ datasets, but trained only on $9177$ user images in the Celeb-A dataset different from the ones tested on. Thus, there is no overlap in identities used in training set of Pipe and those tested.
@@ 6.2 PIPE Against Existing Protection Schemes
Table III shows that Pipe succeeds with a high ASR. When setting up Facial-FE, we configured its parameters to attain $\leq 70\%$ FNR [^5] TPR is 11-FNR. FNR $\leq 70\%$ implies TPR $30\%$. as specified by [2021:facial-fe]. This requires the face vectors to be scaled down. The scales we obtain for Facenet and Arcface embeddings are 5.7 and 1.39 respectively. At this scale, which is required for reasonable accuracy, we notice that the inputs tend to decode to lattice points $c$ that are sparse 512-dimensional vectors with more than 194 zero elements. Consequently, the helper string ss=cxss=c-x contains significant information about $x$ making Facial-FE insecure. This issue cannot be resolved simply by changing the scale factor. As mentioned in [2021:facial-fe] as well, when we experimentally reduce the scale factor to improve security significantly, the TPR becomes unusably small. Thus, the poor trade-off between security and accuracy appears unavoidable for Facial-FE.
@@ 6.3 Prior Attacks vs Partial Protection Schemes
To answer this, we evaluate $3$ state-of-the-art inversion attacks on face recognition models: Bob [2024:bob], KED-MI [2021:ked-mi], GMI [2020:gmi]. Bob is the state-of-the-art white-box embedding inversion attack which has reported better ASR than attacks preceding it. KED-MI and GMI are state-of-the-art black-box model inversion attacks [^6] The core of these attacks only requires black-box access to the target model, and a separate inversion model trained on the public dataset.  [2024:midre]. They both use GANs as the inversion model and have the best-known performance among black-box model inversion attacks [2022:bido].
@@ Results.
Figure 3 presents the representative images reconstructed by Pipe and Bob. Most Pipe reconstructions are accepted as the same face by the target model, which is not the case for Bob. Moreover, as seen in Tables VIII and IX, Pipe achieves a lower (better) visual similarity LPIPS score [2018:LPIPS] and a lower (better) FID score [2017:fid]. This further confirms that Pipe creates perceptually closer reconstructions to the original images.
KED-MI and GMI attacks obtain less than $0.16\%$ ASR on average over 5 trials, with or without a protection scheme. The weak performance of KED-MI and GMI is expected since these attacks rely solely on classification outputs, which are inherently more restrictive than using the embedding vector itself. We qualitatively inspected over $20$ examples returned by these attacks, and we find that the reconstructions are quite far from the original consistently. Our implementation builds on code provided by [2022:bido].
@@ 7 Our Proposed Scheme: L2FE-Hash
The Rep procedure receives $x^{\prime}$. It de-serializes $p$ to obtain $c=Ab+x$ that was computed during Gen. It implicitly checks if $c-x^{\prime}$ decodes to the correct ephemeral secret value by checking if the hash of the decoded value $H_{k}(b)$ matches what is stored in the database. The Decode procedure is the well-known Babai’s Nearest Plane (BNP) algorithm [1986:BNP], which performs the $\ell_{2}$ error correction.
@@
Constructing a fuzzy extractor for an arbitrary family of distributions $\mathcal{X}$ is known to be statistically impossible [2020:fe-impossible]. Therefore, to characterize the computational hardness of breaking L2FE-Hash, one must specialize to specific distributions. If we strictly confine the input face distribution $\cal{X}$ to be discrete Gaussian or uniform over a small interval, then the hardness of recovering $b$ from $(A,Ab+x)$ is exactly the cryptographic hardness of the LWE problem [2005:lwe, 2013:dmq]. Nonetheless, this simplification is too unrealistic, as real-world face distributions are not known to follow any cryptographically hard distribution or have any closed-form mathematical characterization.
As the next best approach, we characterize the support of the distribution $\mathcal{X}$ as a set of balls bounded in some $R,R]^{m}$ space, since our application aims to correct within some $\ell_{2}$ radius for a set of users. We refer to such distributions with high min-entropy $h$ as well-behaved. Under the assumption that $\mathcal{X}$ is uniform over the assumed support, we prove that L2FE-Hash is a secure FE where the number of secure output bits $\kappa$ is upper bounded by some function of $h$.
Theorem 1 makes the claim precise (see Appendix B for the proof). The L2FE-Hash output $r$ is a $\kappa$-bitstring, but equivalently a vector in $\mathbb{Z}_{q}^{n}$ as needed, where $n=\kappa\log_{q}(2)$.
@@
Theorem 1 specifies the minimum input entropy needed for good security for an abstract characterization of the distribution $\cal{X}$. But, determining entropy $h$ for real distributions is intractable. So, we derive practical parameters such that the protected embeddings give good accuracy (TPR over $60\%$) and tractable runtime for Gen and Rep. Note that once TPR exceeds well over $50\%$, majority voting over multiple independent face samples can boost it arbitrarily close to $100\%$ [2011:majority-vote]. To achieve accuracy of $1-\delta$, it suffices to have $O(\log(1/\delta))$ face images of the user at authentication time, which can be run in parallel for efficiency. We evaluate L2FE-Hash using the best-known attacks, including prior ones and our Pipe attack on real datasets, to practically assess security.
@@
We have shown that L2FE-Hash is a provably secure fuzzy extractor for $\ell_{2}$ distance if the input distribution $\cal{X}$ satisfies certain conditions. Next, we state the desired connection between a fuzzy extractor and an ideal defense primitive as Theorem 2. We defer the formal definition of these two primitives and the ideal primitive security game to Appendix A, and the proof of Theorem 2 to Appendix C due to space limits.
@@
Let FE be a ($X,\mathcal{X},\kappa,t,\epsilon_{\text{FE}}$)-fuzzy extractor as in [2004:fe].Then FE is a ($X,\mathcal{X},\{0,1\}^{\kappa},t,\lambda$)-ideal primitive with adversarial advantage $\text{Adv}(\mathcal{A})\leq\epsilon_{\text{FE}}(1-\delta)$.